OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (–1, 5) and has a slope of –4? y = x +

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{slate}{ \displaystyle y-y_{_{^1}}= m\left( x-x_{_{^1}} \right) }\)

OpenStudy (solomonzelman):

Where, \(\large\color{blue}{ \displaystyle \left(x_{_{^1}},y_{_{^1}} \right) }\) is the point and \(\large\color{blue}{ \displaystyle m }\) is the slope.

OpenStudy (anonymous):

so m=-4

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

can you fill in the \(\large\color{slate}{ \displaystyle y-y_{_{^1}}= m\left( x-x_{_{^1}} \right) }\) ?

OpenStudy (anonymous):

is it y=-1x+5

OpenStudy (anonymous):

@satellite73 may you help me please

OpenStudy (solomonzelman):

\(\large\color{slate}{ \displaystyle y-y_{_{^1}}= m\left( x-x_{_{^1}} \right) }\) \(\large\color{slate}{ \displaystyle y-(5)= -4\left( x-(-1)\right) }\)

OpenStudy (solomonzelman):

you need to convert to: \(\large\color{slate}{ y=mx+b }\) that is all.